OpenStudy (anonymous):

cubed root of (-8x^9).... I have -2x^?

OpenStudy (anonymous):



OpenStudy (anonymous):

(-8x^9)^1/3 =-2x^3

OpenStudy (anonymous):

so you just take the square root of the nine

OpenStudy (anonymous):

take the cube root of -8 that is -2 and the x power of 9 divided by 1/3 power=3

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

i mean multiplied by 1/3 or divided by 3

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so what do you do if it says the cubed root of say 2? what is that then?

OpenStudy (anonymous):

ok cube root of 2=2^(1/3) if you use a calculator =1.2599

OpenStudy (anonymous):

ok thanks I'm not really sure how to do this hw I have lol so Im just trying to piece it all together thank you!

OpenStudy (anonymous):